The demand for a good produced by a firm has been reliably measured by P =...

  1. The demand for a good produced by a firm has been reliably measured by P = 100 – 5Q, output Q is measured in thousands of units. If the total cost function is given by: C = 10Q, what is the optimal level of output produced by the monopolist?

MC = 10

MR = 100 – 10Q
Answer #1

Answer : For monopolist the profit-maximizing condition is MR = MC. So,

100 - 10Q = 10

=> 100 - 10 = 10Q

=> 10Q = 90

=> Q = 90 / 10

=> Q = 9

Therefore, here the monopolist's optimal output level is 9 thousands of units.